OpenStudy (anonymous):

compute using polar coordinates to show this integral equals pi {double integral} e^-(x^2+y^2) dx dy, x=-inf..inf, y=-inf..inf

OpenStudy (nikvist):

\[x^2+y^2=\rho^2\quad,\quad dxdy=\rho d\rho d\phi\quad;\quad \rho\in[0,+\infty)\quad,\quad\phi\in[0,2\pi]\]